16 is what percent of 34?
Found 2 solutions by solver91311, Mathtut:
Answer by solver91311(24713)


Divide 16 by 34, then move the decimal point two places to the left.



John


Answer by Mathtut(3670)

16=34x
:
x=16/34=8/17=.4706(100)=47.06%
:
so 16 is 47.06% of 34